# Calibration in Deep Learning: A Survey of the State-of-the-Art

### v3

@@ 5.1 Mitigating Calibration Bias
Accurately and reliably measuring calibration is still challenging due to the introduced biases from the binning mechanism and the finite sample numbers (?). For the former challenge, it mainly suffers from sensitivity and data inefficiency issues. The sensitivity to the binning scheme is presented in 5. We can see that for a given model, increasing bin numbers gives higher ECE and MCE scores. A KDE-based ECE Estimator (?) was proposed to replace histograms with non-parametric density estimators, which are continuous and more data-efficient. Measuring the bias is then important for having a correct calibration evaluation. Roelofs et al. (?) proposed Bias-by-Construction (BBC) to model the bias in bin-based ECE as a function of the number of samples and bins. It confirms the existence of non-negligible statistical biases. To follow this line of work, future efforts will include developing an unbiased calibration estimator, exploring the trade-off between calibration bias and variance as mentioned in  (?).
@@ 5.2 Calibrating Generative Models
Recent efforts in model calibration have predominantly centered around classification and regression tasks, leaving a noticeable gap in discussions regarding calibration for sequence generation. While advancements in calibration techniques have greatly benefited these tasks, the applicability and efficacy of such techniques in sequence generation remain under-explored. Building upon the seminal work of  (?), which highlighted the inadequately calibrated token-level probabilities in neural machine translation tasks, there is a pressing need to extend these insights to a broader range of generative models. The findings of (?, ?) revealed a crucial connection between token-level probability miscalibration and the counter-intuitive drop in BLEU scores with increased beam size, as noted by (?). Furthermore, they demonstrated that improved sequence-level calibration can be achieved through the re-calibration of token-level probabilities. This underscores the importance of addressing token-level probability miscalibration not only for enhancing performance metrics but also for improving the overall reliability and coherence of generated sequences.
Expanding upon these observations, recent studies such as that by (?) have confirmed the presence of token-level probability miscalibration in Large Language Models (LLMs) utilized for few-shot learning tasks. The identified culprit, the softmax bottleneck as elucidated by (?), particularly on a large vocabulary, further exacerbates this issue. In the context of sequence generation, early token probability miscalibration can have cascading effects, magnifying errors throughout the entire sequence.
Moving forward, addressing the challenge of effectively calibrating token-level probabilities in various settings represents a promising avenue for research, especially in the era of LLMs. Extending the findings from classification and regression tasks to sequence generation domains requires innovative approaches tailored to the unique characteristics and challenges posed by generative models. Insights gained from recent advancements such as generative calibration (?), sequence likelihood calibration (?) and its extension (?).

### v4

@@ 7.1 Mitigating Calibration Bias
Accurately and reliably measuring calibration is still challenging due to the introduced biases from the binning mechanism and the finite sample numbers (?). For the former challenge, it mainly suffers from sensitivity and data inefficiency issues. The sensitivity to the binning scheme is presented in Figure 4. We can see that for a given model, increasing bin numbers gives higher ECE and MCE scores. A KDE-based ECE Estimator (?) was proposed to replace histograms with non-parametric density estimators, which are continuous and more data-efficient. Measuring the bias is then important for having a correct calibration evaluation. Roelofs et al. (?) proposed Bias-by-Construction (BBC) to model the bias in bin-based ECE as a function of the number of samples and bins. It confirms the existence of non-negligible statistical biases. To follow this line of work, future efforts will include developing an unbiased calibration estimator, exploring the trade-off between calibration bias and variance as mentioned in  (?).
@@ 7.2 Calibrating Generative Models
Recent efforts in model calibration have predominantly centered around classification and regression tasks, leaving a noticeable gap in discussions regarding calibration for sequence generation. While advancements in calibration techniques have greatly benefited these tasks, the applicability and efficacy of such techniques in sequence generation remain under-explored. Building upon the seminal work of  (?), which highlighted the inadequately calibrated token-level probabilities in neural machine translation tasks, there is a pressing need to extend these insights to a broader range of generative models. The findings of (?, ?) revealed a crucial connection between token-level probability miscalibration and the counter-intuitive drop in BLEU scores with increased beam size, as noted in (?). Furthermore, they demonstrated that improved sequence-level calibration can be achieved through the re-calibration of token-level probabilities. This underscores the importance of addressing token-level probability miscalibration not only for enhancing performance metrics but also for improving the overall reliability and coherence of generated sequences.
Building on these insights, recent studies such as Zhao et al. (?) have demonstrated that token-level probability miscalibration is a persistent issue in Large Language Models (LLMs), particularly in few-shot learning scenarios. One key contributor is the softmax bottleneck—as identified in (?)—which becomes especially problematic in models with large vocabularies. In sequence generation tasks, early-stage miscalibrations can compound across tokens, leading to amplified errors throughout the generated output.
Looking ahead, developing methods to accurately calibrate token-level probabilities in generative settings is an important and emerging direction, particularly in the context of LLMs. Unlike classification or regression, sequence generation introduces unique challenges that demand novel calibration techniques. Recent efforts such as generative calibration (?), sequence likelihood calibration (?), and its enhanced variant SLIC (?) offer promising starting points for advancing calibration in this space.